# Geometry Aware Operator Transformer As An Efficient And Accurate Neural Surrogate For PDEs On Arbitrary Domains

### v2

@@ D.1 Poisson-C-Sines
where $r=-0.5$ and $K=16$. The coefficients $a_{ij}$ are sampled i.i.d. uniformly from $[-1,1]$ to generate the dataset. The solution $u$ is computed on a circular domain with zero Dirichlet boundary conditions. The dataset is generated using a finite element method (FEM) on a triangular mesh in a circular domain. The mesh is generated using the Delaunay algorithm with 16431 points and 32441 elements.
@@ D.2 Compressible Flow Past Airfoils & Bluff Bodies
A classic benchmark for compressible flow physics used for testing the accuracy of neural operators and PDE foundation models is the case of flow past airfoils [20, 26]. The datasets used in these papers are limited to transonic flow past perturbations of a single airfoil. To capture a broader range of rich flow phenomena, it is essential to explore the parameter space spanned by the Mach number $Ma$, angle of attack $\alpha$ and the shape function. To address this issue, this new dataset introduces samples comprising a range of flow phenomena from subsonic to supersonic flow for varying angles of attack across classical airfoils and various bluff body geometries. The steady-state compressible Euler equations govern the flow phenomena in this dataset. The equations have been solved using the finite-volume EULER solver of the open-source software SU2 [11] on an unstructured grid generated by Delaundo [40]. Convective flux discretization is done using the Jameson-Schmidt-Turkel (JST) scheme that is designed especially for achieving quick convergence to steady-state solutions of the compressible Euler equations. Figure D.1 represents an O-type unstructured mesh generated using Delaundo for the RAE2822 airfoil. Similar O-type unstructured meshes have been generated for all airfoils and bluff-bodies considered. The free-stream pressure and temperature conditions for all simulations in this dataset are $p_{\infty}=1\text{ atm and }T_{\infty}=288.15\text{ K}$.
@@ D.2.1 Airfoils
Figure D.2: Adaptively refined meshes for flow past the RAE2822 airfoil at $\alpha=2.0$ and at different $Ma$ = 0.8 (left), 1.0 (center), 1.4 (right).
@@
$S^{\text{U}}(\xi)=\sum_{i=0}^{n}A_{i}\frac{n!}{i!(n-i)!}\xi^{i}(1-\xi)^{i},\quad S^{\text{L}}(\xi)=\sum_{i=0}^{n}B_{i}\frac{n!}{i!(n-i)!}\xi^{i}(1-\xi)^{i}$ (D.8)
The polynomials $S_{i,n}=\frac{n!}{i!(n-i)!}\xi^{i}(1-\xi)^{i}$ associated with the coefficients ($A_{i},B_{i}$) are Bernstein polynomials and $n=7$ is chosen. The parameters ($A_{i},B_{i}$) directly influence key airfoil design variables such as the leading edge radius, trailing edge boattail angle, maximum airfoil thickness and maximum thickness location. The parameters $(A_{0},B_{0})$ are linked to the leading edge radius $R_{\text{LE}}$ as follows,
@@ D.2.2 Bluff-Body
We sample 4384 solutions from our dataset with a train/validation/test split of 4000/128/256 for all bluff-body geometries used for "Training and Testing" . For each data, we sub-samples 14000 points for training, validation and testing.
@@ Appendix E Additional Results
E.1 Accuracy, Robustness and Computational Efficiency Metrics
@@
For the pressure and wall shear stress (WSS) in the DrivAerNet++ dataset, we evaluated the model on 1154 samples according to the official leaderboard. The errors are calculated based on normalized pressure and WSS. For pressure, the mean and standard deviation (std) for normalization were obtained from the open-sourced code of [12]. However, for WSS, as the normalization statistics were not open-sourced, we calculated the mean and variance for the x, y, and z components over 8000 samples to be used for normalization. The Mean Squared Error (MSE) and Mean Absolute Error (MAE) are first computed for each individual sample. Then, the average of these errors across the 1154 test samples is reported as the final result. This entire procedure strictly follows their open-sourced code methodology of [12].
@@
Robustness $\text{Robustness Score}=\bar{S}_{\text{norm}}\times(1-\text{CV}),$ (E.3)
@@
E.2 Results for Radar Chart in Main Text.
@@
E.3 Regular Grid Dataset
@@
Structured GAOT RIGNO-18 RIGNO-12 CNO scOT FNO
NS-Gauss 2.29 2.74 3.78 10.9 2.92 14.41
NS-PwC 1.23 1.12 1.82 5.03 7.12 12.55
NS-SL 0.98 1.13 1.82 2.12 2.49 2.08
NS-SVS 0.46 0.56 0.75 0.70 1.01 7.52
CE-Gauss 5.28 5.47 7.56 22.0 9.44 28.69
CE-RP 4.98 3.49 4.43 18.4 9.74 38.48
Wave-Layer 5.40 6.75 8.97 8.28 13.44 28.13
E.4 Model and Dataset Scaling
@@
To further investigate the scalability of our approach, we conduct an ablation study on how different model sizes affect performance. We focus on the two compressible Euler datasets, CE-Gauss and CE-RP, each with $1{,}024$ training trajectories. We measure the final-time relative $L^{1}$ error ($t=t_{14}$) and record the total number of parameters and per-epoch training time under various hyperparameter configurations.
@@
[…]
@@
[…]
@@ E.5 Resolution Invariance
E.6 Transfer Learning
@@ F.2 Tokenization Strategies
Results show that fewer tokens (e.g., [32,32]) can degrade performance in some cases (elasticity), presumably because the domain coverage becomes coarser, making it harder to capture local variations. More tokens ([64,64] or [128,128]) typically improve accuracy and stabilize convergence. Nevertheless, computational costs rise when the number of tokens grows, as transformer attention scales quadratically with token count. Increasing the patch size (PS) reduces the number of tokens entering the transformer, lowering the training time. Encouragingly, performance does not degrade sharply with larger patches. For instance, going from PS $=2$ to $8$ is fairly stable across datasets. Note that the overall parameter count can increase if each token aggregates larger local features, but in practice, training runs faster due to fewer tokens in self-attention. Radius (GR) grows if we reduce the number of latent tokens because we need to ensure coverage of the entire physical domain by enlarging the receptive field. This is critical for unstructured or irregular samples, especially if tokens must capture a bigger subregion.

### v3

@@ D.1 Poisson-C-Sines
where r=0.5r=-0.5 and $K=16$. The coefficients $a_{ij}$ are sampled i.i.d. uniformly from [1,1][-1,1] to generate the dataset. The solution $u$ is computed on a circular domain with zero Dirichlet boundary conditions. The dataset is generated using a finite element method (FEM) on a triangular mesh in a circular domain. The mesh is generated using the Delaunay algorithm with 16431 points and 32441 elements.
@@ D.2 Compressible Flow Past Airfoils & Bluff Bodies
A classical benchmark for compressible flow physics used for testing the accuracy of neural operators and PDE foundation models is the case of flow past airfoils [21, 25]. The datasets used in these papers are limited to transonic flow past perturbations of a single airfoil. To capture a broader range of rich flow phenomena, it is essential to explore the parameter space spanned by the Mach number $Ma$, angle of attack $\alpha$ and the shape function. To address this issue, this new dataset introduces samples comprising a range of flow phenomena from subsonic to supersonic flow for varying angles of attack across classical airfoils and various bluff body geometries. The steady-state compressible Euler equations govern the flow phenomena in this dataset. The equations have been solved using the finite-volume EULER solver of the open-source software SU2 [12] on an unstructured grid generated by Delaundo [40]. Convective flux discretization is done using the Jameson-Schmidt-Turkel (JST) scheme that is designed especially for achieving quick convergence to steady-state solutions of the compressible Euler equations. Figure D.1 represents an O-type unstructured mesh generated using Delaundo for the RAE2822 airfoil. Similar O-type unstructured meshes have been generated for all airfoils and bluff-bodies considered. The free-stream pressure and temperature conditions for all simulations in this dataset are $p_{\infty}=1\text{ atm and }T_{\infty}=288.15\text{ K}$.
@@ D.2.1 Airfoils
Figure D.2: Adaptively refined meshes for flow past the RAE2822 airfoil at $\alpha=2.0$ and at different MaMa = 0.8 (left), 1.0 (center), 1.4 (right).
@@
$S^{\text{U}}(\xi)=\sum_{i=0}^{n}A_{i}\frac{n!}{i!(n-i)!}\xi^{i}(1-\xi)^{n-i},\quad S^{\text{L}}(\xi)=\sum_{i=0}^{n}B_{i}\frac{n!}{i!(n-i)!}\xi^{i}(1-\xi)^{n-i}$ (D.8)
The polynomials $S_{i,n}=\frac{n!}{i!(n-i)!}\xi^{i}(1-\xi)^{n-i}$ associated with the coefficients ($A_{i},B_{i}$) are Bernstein polynomials and $n=7$ is chosen. The parameters ($A_{i},B_{i}$) directly influence key airfoil design variables such as the leading edge radius, trailing edge boattail angle, maximum airfoil thickness and maximum thickness location. The parameters $(A_{0},B_{0})$ are linked to the leading edge radius $R_{\text{LE}}$ as follows,
@@ D.2.2 Bluff-Body
We sample 4384 solutions from our dataset with a train/validation/test split of 4000/128/256 for all bluff-body geometries used for "Training and Testing" . For each data, we sub-sample 14000 points for training, validation and testing.
@@ Appendix E Additional Results
E.1 Asymptotic Complexity
Table E.1 summarizes the asymptotic complexities of the compared models in Fig. 1. In practice, the number of edges $E$ is typically $3-5\times$ the number of nodes $N$. For consistency, we denote the number of latent tokens by $T$ across all models; however, ViT-style patchifying in GAOT effectively reduces $T$ and thereby improves scalability. Furthermore, different models employ distinct hidden dimensions $C$ for their edge update functions, which introduces further variation in the total computational cost.
Table E.1: Asymptotic complexity. NN: number of nodes, $E$: number of edges, $T$: latent tokens, $C$: hidden channels, $p$: patch size, $L$: number of layers/blocks.
Model Total Complexity
GAOT $O(E+L(T/p)^{2}C)$
GINO $O(E+LT\log T)C$
RIGNO $O(LEC)$
Transolver $O\big(L(NTC+T^{2}C)\big)$
The main computational burden in graph-based models arises from message passing, and memory usage is dominated by storing edge activations, scaling linearly with both $E$ and $C$. GAOT, RIGNO, and GINO compress information into latent tokens and then perform $L$ layers of latent processing. In contrast, Transolver performs compression and decoding within each block, operating directly on nodes of $N$ rather than edge $E$, but accumulating higher total computation across layers.
It is worth noting that these are theoretical costs only. On modern hardware, actual runtime can deviate significantly because graph-based operations rely on sparse and irregular memory access patterns, which are inefficient on GPUs. By contrast, transformer-style dense computations are highly optimized and throughput-efficient on current accelerators. The following Sec E.2 provides a more practical profiling comparison.
E.2 Runtime Profiling
To further analyze the computational efficiency of our implementation, we compare the detailed runtime breakdowns of GINO [28] and GAOT on the NACA0012 dataset using a single NVIDIA GeForce RTX 4090 GPU. Table E.2 presents the per-component execution times for both models at a batch size of 1.
Table E.2: Comparison of time breakdown between GINO (6.07M parameters) and GAOT (5.62M parameters) with batch size = 1 on the NACA0012 dataset.
Component GINO GAOT
Sub-Operation Time (ms) Sub-Operation Time (ms)
Encoder Graph building 1.984 AGNO 2.296
GNO 2.793 GeoEmb 3.412
Total Encoder 4.777 Total Encoder 5.708
Processor Per-layer FNO 1.178 Per-layer Transformer (ps=2) 2.02
$1.178\times 5$ layers 5.89 $2.02\times 5$ layers 10.1
Decoder Graph building 1.984 AGNO 1.356
GNO 1.140 GeoEmb 1.865
Total Decoder 3.124 Total Decoder 3.221
Total Time 13.791 19.029
For GINO, the total runtime per sample is 13.79 ms, where the encoder and decoder each spend almost 2 ms on graph building, comparable to or even exceeding the actual GNO operations. Also, the official implementation of GINO does not support batch processing when the underlying graphs change within a batch, as is typical for PDEs defined on arbitrary domains. This limitation arises because batch-processing the encoder and decoder often leads to out-of-memory (OOM) errors, making GINO scale poorly with batch size.
In contrast, GAOT eliminates the runtime overhead of graph building by precomputing graphs and loading them efficiently from cache. More importantly, to overcome the memory bottleneck in batch processing, we adopt a hybrid execution strategy: the encoder and decoder are processed sequentially through a lightweight for-loop, which avoids memory overflow without introducing noticeable computational overhead, while the processor (Transformer) is fully batch-processed to maximize throughput. This design choice is based on the observation that, for larger models, the processor typically dominates the compute cost, making it the optimal module to parallelize. As shown in Table E.3, GAOT achieves a total runtime of 263.1 ms for batch size 32, while GINO, lacking batching support, must loop 32 times, resulting in $13.791\times 32=441.3$ ms. The Transformer processor scales efficiently (from 10.1 ms for batch size 1 to only 30.1 ms for batch size 32), while maintaining a manageable total memory usage of 7.74 GB. Consequently, GAOT achieves significantly better scalability than GINO in both input size and model size, without sacrificing computational efficiency. For very large graphs, e.g. DrivAerML dataset [2], edge masking is used to control memory usage ensuring stable training even on limited-memory GPUs.
Table E.3: Breakdown of time and memory for GAOT (5.62M parameters) with batch size = 32 on the NACA0012 dataset.
Component Sub-Operation Time (ms) Memory (GB)
Encoder AGNO + GeoEmb 126 1.836
Processor 5-layer Transformer, ps=2 30.1 3.592
Decoder AGNO + GeoEmb 107 2.310
Total 263.1 7.738
E.3 Accuracy, Robustness and Computational Efficiency Metrics
@@
For the pressure and wall shear stress (WSS) in the DrivAerNet++ dataset, we evaluated the model on 1154 samples according to the official leaderboard. The errors are calculated based on normalized pressure and WSS. For pressure, the mean and standard deviation (std) for normalization were obtained from the open-source code of [13]. However, for WSS, as the normalization statistics were not open-sourced, we calculated the mean and variance for the x, y, and z components over 8000 samples to be used for normalization. The Mean Squared Error (MSE) and Mean Absolute Error (MAE) are first computed for each individual sample. Then, the average of these errors across the 1154 test samples is reported as the final result. This entire procedure strictly follows their open-source code methodology of [13].
@@
$\text{Robustness Score}=\bar{S}_{\text{norm}}\times(1-\text{CV}),$ (E.3)
@@
E.4 Results for Radar Chart in Main Text.
@@
E.5 Regular Grid Dataset
@@
Furthermore, we ablated GAOT further by just running the ViT processor (without the encoder/decoder) on these Cartesian datasets. From these results, it is clear that in some cases, ViT is comparable to or slightly better than GAOT, indicating that the power of GAOT in these cases stemmed from a good processor. However in many more cases, GAOT is significantly superior to ViT also showing that the encoder/decoder contribute significantly to expressivity and together, this combination achieves SOTA performance. This interesting experiment clearly delineates the relative contributions of encoder/decoder vs. processor.
@@
Structured GAOT RIGNO-18 RIGNO-12 CNO ViT scOT FNO
NS-Gauss 2.29 2.74 3.78 10.9 3.16 2.92 14.41
NS-PwC 1.23 1.12 1.82 5.03 3.89 7.12 12.55
NS-SL 0.98 1.13 1.82 2.12 0.73 2.49 2.08
NS-SVS 0.46 0.56 0.75 0.70 0.39 1.01 7.52
CE-Gauss 5.28 5.47 7.56 22.0 6.81 9.44 28.69
CE-RP 4.98 3.49 4.43 18.4 4.30 9.74 38.48
Wave-Layer 5.40 6.75 8.97 8.28 5.48 13.44 28.13
E.6 Model and Dataset Scaling
@@
To further investigate the scalability of our approach, we conduct an ablation study on how different model sizes affect performance. We focus on the two compressible Euler datasets, CE-Gauss and CE-RP, each with 1,0241{,}024 training trajectories. We measure the final-time relative $L^{1}$ error ($t=t_{14}$) and record the total number of parameters and per-epoch training time under various hyperparameter configurations.
@@
Finally, the bottom block (rows 8–10) examines the effect of transformer layers TL from $1$ up to $10$. With TL=1, errors remain quite high ($25.0\%$ on CE-Gauss); adding layers substantially reduces error to $9.0\%$ at $TL=3$, and ultimately down to $5.28\%$ at TL=10 on CE-Gauss. Increasing TL to 10 also expands the parameter count to $11.2\text{M}$, nearly doubling the training time per epoch (260s).
Figure E.2 illustrates how errors decrease as we scale THS or TL, while Figure E.3 shows the effect of changing LC or the total parameter count. The largest model tested reaches $11.2\text{M}$ parameters and attains around $5\%$ error on CE-Gauss, demonstrating the potential to improve accuracy by investing in more computational resources.
@@
Time-Dependent (Fluid) Datasets. We plot the final-time ($t=t_{14}$) error on multiple fluid PDE benchmarks as a function of the training set size $\{128,\,256,\,512,\,1024\}$. All of these datasets use partial-grid subsampling. The results confirm that as we increase the number of training trajectories, errors consistently drop across all fluid datasets, often in a near-linear fashion with respect to the number of training trajectories. This trend highlights the model’s capacity to benefit from additional time-series diversity.
Time-Independent (Static) Datasets. We similarly measure how the final solution error decreases when expanding the dataset size to $\{128,\,256,\,512,\,1024,\,2048\}$ for three static PDE tasks: Poisson-Gauss, Poisson-C-Sines, and Elasticity. Note that elasticity is limited to at most $1024$ samples due to data availability. Across all these static problems, we observe a consistent downward slope in error as the number of samples increases, again underscoring the advantage of larger training sets.
@@ E.7 Resolution Invariance
E.8 Comparison of Neural Field with UPT
The neural field property enables continuous evaluation of the learned mapping at arbitrary spatial points, allowing models to train and test at different resolutions-a highly desirable capability for PDE operator learning [1, 48]. This feature allows a model trained on a coarser resolution for efficiency to be tested at finer resolutions for higher accuracy.
In Table 3 of the main paper, we demonstrated this property for GAOT on the industrial-scale 3D Drivaernet++ dataset. However, as noted in this result, possessing the neural field property does not necessarily imply higher accuracy. Indeed, GAOT trained using only 10% of input points but tested on the full resolution (500K surface points) is less accurate than the fully trained version, though still outperforming several baselines.
To further examine this, we compared GAOT with the neural-field-based UPT model [1] on the 2D NACA0012 dataset under two settings: (i) both trained and tested at full resolution, and (ii) trained at one-fourth resolution but tested at full resolution. The results are shown in Table E.7.
Table E.7: Neural field training with 1/4 sub-sampling on the NACA0012 dataset.
Model Error (Full Resolution) Error (Neural Field)
GAOT 6.81 9.54
UPT 16.1 16.3
These results reveal that: (a) training at full resolution yields better accuracy than coarse training; (b) GAOT consistently outperforms UPT under both configurations; and (c) UPT’s limited improvement with resolution suggests poor convergence, further highlighting the robustness of GAOT as a neural surrogate with flexible spatial generalization.
E.9 Transfer Learning
@@ E.10 Training Randomness
E.11 Runtime Comparison between GAOT and Classical Solvers
The main rationale for the design of efficient neural surrogates such as GAOT lies in the fact that classical numerical PDE solvers are slow, particularly in industrial 3D problems. On the other hand, neural operators are ultra-fast to infer. We discuss this issue from the perspective of GAOT in this section. To begin with, Table E.9 summarizes the inference times of GAOT and compares them with the runtimes of classical numerical solvers across several CFD datasets. For fairness, all GAOT inference times include graph construction overhead.
Table E.9: Comparison of GAOT inference time with traditional CFD solvers.
Dataset Mesh Points Inference Time (ms) Traditional Solver Speedup
Bluff-body 14,000 9.77 $283\sim 419$ s $\sim 3\times 10^{4}$
NS-SL 16,384 10.14 0.1 s $\sim 10^{1}$
DrivaerNet++ 500,000 365.36 375 core hours $\sim 3.7\times 10^{7}$
DrivaerML 9,000,000 14,091.08 61,440 core-hours $\sim 1.5\times 10^{9}$
The NS-SL dataset is taken from Ref. [21], following the setup of Ref. [39] with random point-cloud inputs. Their optimized GPU-based spectral viscosity solver achieves approximately 0.1s per sample. For our new datasets-NACA0012, NACA2412, RAE2822, and Bluff-body-which simulate compressible flow around airfoils and bluff bodies, the traditional CPU-based solver required between 283s and 419s per sample due to adaptive grids and iterative refinement (see SM D for details). For the large-scale 3D benchmark DrivaerNet++ [13], the original authors reported an average cost of approximately 375 CPU hours per sample. For DrivAerML, [2] indicates that each scale-resolving CFD simulation using the hybrid RANS–LES approach required around 40 hours on 1536 cores ($\approx 61,440$ core-hours) on AWS HPC clusters, which represents an industrial-grade high-fidelity CFD workflow that includes near-wall resolution and statistical convergence monitoring.
Overall, traditional numerical solvers span a runtime range from 0.1s to a few minutes for 2D problems discussed in this work, and up to tens of thousands of core-hours for industrial 3D simulations. In contrast, GAOT performs inference within 8.95-10.14 ms for 2D datasets and around 365ms for 3D DrivaerNet++, achieving a remarkable speedup of 1–5 orders of magnitude for 2D problems and up to 10 orders of magnitude for 3D industrial-scale benchmarks. The reason why DrivAerML shows a significantly higher runtime compared to other datasets is mainly due to the excessive time spent on graph building. In fact, if we exclude the graph building process, its inference speed is 446.06 ms. The relatively long graph building time can be attributed to two factors: first, the large scale of the input mesh points; and second, to control memory usage while covering the entire solution domain, we employed more complex graph building techniques (see Section B.5). Nevertheless, even under these conditions, the efficiency improvement compared to traditional solvers remains at the billion order of magnitude level as its runtime of approximately $15$ secs is completely dwarfed by the more than the $60K$ core hours run time of the LES based ground truth simulations.
One can argue that the runtime of classical numerical PDE solvers can be reduced by coarsening the resolution. This does not hold for the industrial scale datasets as there is a minimum mesh resolution that is essential for the underlying physics to be resolved. However, on the academic 2D datasets, we can perform such a coarsening and examine if the cost-accuracy pay-off for GAOT still holds. To this end, we consider the NACA0012 dataset and coarsen the resolution of the underlying finite-volume solver by two levels of refinement. The underlying error is now $4.5\%$, when compared to the ground truth but the run-time reduces from $7$ mins to approximately $1$ minute. This error is comparable to the error of GAOT ($6.8\%$), while the runtime of GAOT for this dataset is approximately $10$ milli-seconds, leading to a speedup of $6000$. This further demonstrates the enormous advantage that an efficient and accurate neural operator such as GAOT can provide, when compared to classical numerical PDE solvers.
@@ F.2 Tokenization Strategies
Results show that fewer tokens (e.g., [32,32]) can degrade performance in some cases (elasticity), presumably because the domain coverage becomes coarser, making it harder to capture local variations. More tokens ([64,64] or [128,128]) typically improve accuracy and stabilize convergence. Nevertheless, computational costs rise when the number of tokens grows, as transformer attention scales quadratically with token count. Increasing the patch size (PS) reduces the number of tokens entering the transformer, lowering the training time. Encouragingly, performance does not degrade sharply with larger patches. For instance, going from PS =2=2 to $8$ is fairly stable across datasets. Note that the overall parameter count can increase if each token aggregates larger local features, but in practice, training runs faster due to fewer tokens in self-attention. Radius (GR) grows if we reduce the number of latent tokens because we need to ensure coverage of the entire physical domain by enlarging the receptive field. This is critical for unstructured or irregular samples, especially if tokens must capture a bigger subregion.
@@ F.4 Geometric Embedding
We further conducted ablation studies on DrivaerNet++ and DrivaerML to examine the effect of statistical geometric embeddings. Three configurations were considered: (1) without geometric embedding (nGEmb-nGEmb), (2) applying geometric embedding only in the encoder (GEmb-nGEmb), and (3) applying geometric embedding in both the encoder and decoder (GEmb-GEmb).
Table F.5: Ablation study on the effect of geometric embedding configurations across DrivAerNet++ and DrivAerML datasets.
Dataset GEmb-nGEmb GEmb-GEmb nGEmb-nGEmb
MSE Mean AE MSE Mean AE MSE Mean AE
DrivAerNet++(p) 4.2694 1.0699 4.3119 1.0818 4.5278 1.1036
DrivAerNet++(wss) 8.6878 1.5429 8.7783 1.5690 9.3192 1.6125
DrivAerML(p) 5.1729 1.2352 10.8591 1.7344 8.6625 1.5693
DrivAerML(wss) 16.9818 2.1640 41.0027 3.0822 24.8614 2.4965
The results show that introducing geometric embeddings consistently improves model performance compared to not using them. Interestingly, we observed that applying the geometric embedding only in the encoder yields even better performance, but less computational effort. Moreover, for the DrivAerML dataset, adding a geometric embedding in the decoder actually degrades performance, producing results even worse than the model without any geometric embedding. Our analysis indicates that while the model’s training loss continues to decrease, its validation loss quickly saturates. We think that incorporating geometric embedding in the decoder, where operations are very close to the final predictions, makes the model more prone to overfitting. Especially on the 3D industrial dataset, where we used more complex graph-building techniques. These complex graph-building tricks lack the unified principles for the model to learn and likely amplify the overfitting tendency when the decoder also encodes geometric information. Therefore, we only use geometric embeddings in the encoder for our 3D industrial benchmarks.